OpenStudy (mrmanman305):

cot ( x - pi/2)= -tan x Can anybody help verify???? Medal... @phi @mathstudent55 @mathmale

OpenStudy (phi):

do you know the formula for the tangent of the difference of two angles ?

OpenStudy (anonymous):

\[\cot \left( x-\frac{ \pi }{ 2} \right)=\cot \left\{ -\left( \frac{ \pi }{ 2 } -x\right) \right\}=-\cot \left( \frac{ \pi }{ 2 }-x \right)=-\tan x\]

OpenStudy (mrmanman305):

Is it cot(pi/2-theta)=tan

OpenStudy (mrmanman305):

I didnt know you were able to just multiply by -1

OpenStudy (mrmanman305):

Is that all you have to do?

OpenStudy (anonymous):

no i have not multiplied by -1 if you do not follow i do otherwise. \[\cot \left( -\theta \right) =-\cot \theta \]

OpenStudy (mrmanman305):

ohhhh I get it so you're basically just manipulating the negative sign

OpenStudy (mrmanman305):

Thank you